# Extragradient methods with complexity guarantees for hierarchical variational inequalities

### v1

@@ 4.1 Energetic bounds
$=\varphi^{({\mathsf{F}}_{2},g_{2})}(z^{\ast},\sigma_{k}p^{\ast})-\operatorname{\mathtt{s}}(\sigma_{k}p^{\ast}|\mathcal{S}_{2})$
$\leq\sup_{z\in\mathcal{S}_{2}}\varphi^{({\mathsf{F}}_{2},g_{2})}(z,\sigma_{k}p^{\ast})-\operatorname{\mathtt{s}}(\sigma_{k}p^{\ast}|\mathcal{S}_{2}).$
We note that (A.2) in Appendix A guarantees that the upper bound is informative, in the sense that
$\sup_{z\in\mathcal{S}_{2}}\varphi^{({\mathsf{F}}_{2},g_{2})}(z,\sigma_{k}p^{\ast})-\operatorname{\mathtt{s}}(\sigma_{k}p^{\ast}|\mathcal{S}_{2})\geq 0.$
Plugging this back into (4.15), we obtain
$E_{k+1}(z^{\ast})+D_{k+1}$ $\leq E_{k}(z^{\ast})+D_{k}-\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2}+t_{k}\left[\sup_{z\in\mathcal{S}_{2}}\varphi^{({\mathsf{F}}_{2},g_{2})}(z,\sigma_{k}p^{\ast})-\operatorname{\mathtt{s}}(\sigma_{k}p^{\ast}|\mathcal{S}_{2})\right].$ (4.16)
We see from these estimates that the summability condition (3.4) gives a bound on the sequence $E_{k}(z^{\ast})+D_{k}$. Defining $W_{k}(z^{\ast})\triangleq E_{k}(z^{\ast})+D_{k}$, and
$a_{k}\triangleq t_{k}\left[\sup_{z\in\mathcal{S}_{2}}\varphi^{({\mathsf{F}}_{2},g_{2})}(z,\sigma_{k}p^{\ast})-\operatorname{\mathtt{s}}(\sigma_{k}p^{\ast}|\mathcal{S}_{2})\right],$
we obtain from (4.16)
$W_{k+1}(z^{\ast})-W_{k}(z^{\ast})+\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2}\leq a_{k}.$
From this, summability of $a_{k}$, and non-negativity of $W_{k}(z^{*})$ it follows from Lemma 2.1 that $\lim_{k\to\infty}W_{k}(z^{\ast})=W_{\infty}(z^{\ast})$ exists, and $\lim_{k\rightarrow\infty}\norm{z^{k+1/2}-z^{k}}=0$. By the definition of $E_{k}(z^{*})$, we further obtain that the sequences $(z^{k})_{k\geq 1}$ and $(z^{k+1/2})_{k\geq 1}$ are bounded. $\blacksquare$
@@
We next consider a compact convex set $\mathcal{U}_{2}\subset\operatorname{dom}g_{1}\cap\operatorname{dom}g_{2}$ and choose as anchor point an arbitrary element $z\in\mathcal{U}_{2}$. Then, we have
$\langle{\mathsf{F}}_{1}(z^{k+1/2}),z-z^{k+1/2}\rangle$ $\leq\norm{{\mathsf{F}}_{1}(z^{k+1/2})}\cdot\norm{z-z^{k+1/2}}$ (4.17)
$\leq\left(\norm{{\mathsf{F}}_{1}(z^{1})}+\norm{{\mathsf{F}}_{1}(z^{k+1/2})-{\mathsf{F}}_{1}(z^{1})}\right)\cdot\left(\norm{z^{1}-z^{k+1/2}}+\norm{z^{1}}+\norm{z}\right)$ (4.18)
$\leq\left(\norm{{\mathsf{F}}_{1}(z^{1})}+L_{{\mathsf{F}}_{1}}R\right)\cdot(R+\norm{z^{1}}+\max_{z\in\mathcal{U}_{2}}\norm{z}),$ (4.19)
@@
$C_{\mathcal{U}_{2}}$ $\triangleq\left(\norm{{\mathsf{F}}_{1}(z^{1})}+L_{{\mathsf{F}}_{1}}R\right)(R+\norm{z^{1}}+\max_{z\in\mathcal{U}_{2}}\norm{z})+\Omega_{1}(\mathcal{U}_{2},R)$
to obtain
$\langle{\mathsf{F}}_{1}(z^{k+1/2}),z-z^{k+1/2}\rangle+g_{1}(z)-g_{1}(z^{k+1/2})\leq C_{\mathcal{U}_{2}},\qquad\forall z\in\mathcal{U}_{2},k\geq 1.$
It follows by the definition of $\Psi_{k}(z)$ and monotonicity of ${\mathsf{F}}_{2}$ that, for all $z\in\mathcal{U}_{2}$ and $k\geq 1$,
$\begin{split}\Psi_{k}(z)+\sigma_{k}C_{\mathcal{U}_{2}}&\geq\Psi_{k}(z)+\sigma_{k}(\langle{\mathsf{F}}_{1}(z^{k+1/2}),z-z^{k+1/2}\rangle+g_{1}(z)-g_{1}(z^{k+1/2}))\\ &=\langle{\mathsf{F}}_{2}(z^{k+1/2}),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z)\\ &\geq\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z).\end{split}$ (4.20)
Multiplying by $t_{k}$ and using (4.12), where we now neglect the non-positive term $\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2}$, we obtain, for all $z\in\mathcal{U}_{2},k\geq 1$,
$t_{k}(\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z))$ $\leq t_{k}\Psi_{k}(z)+t_{k}\sigma_{k}C_{\mathcal{U}_{2}}$
$\leq(E_{k}(z)+D_{k})-(E_{k+1}(z)+D_{k+1})+t_{k}\sigma_{k}C_{\mathcal{U}_{2}}.$ (4.21)
Summing (4.21) for $k=1,\ldots,K$ and dividing by $T_{K}$ yields, for all $z\in\mathcal{U}_{2}$,
$\frac{1}{T_{K}}\sum_{k=1}^{K}t_{k}\left(\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z)\right)\leq\frac{1}{T_{K}}(E_{1}(z)+D_{1})+\frac{\sum_{k=1}^{K}t_{k}\sigma_{k}}{T_{K}}C_{\mathcal{U}_{2}}.$ (4.22)
@@
$\langle{\mathsf{F}}_{2}(z),\bar{z}^{K}-z\rangle+g_{2}(\bar{z}^{K})-g_{2}(z)\leq\frac{1}{T_{K}}(E_{1}(z)+D_{1})+\frac{\sum_{k=1}^{K}t_{k}\sigma_{k}}{T_{K}}C_{\mathcal{U}_{2}}\hskip 17.00024pt\forall z\in\mathcal{U}_{2}.$
Thus, for any compact subset $\mathcal{U}_{2}\subset\operatorname{dom}g_{1}\cap\operatorname{dom}g_{2}\subset\mathcal{Z}$ we can use the definition of the localized gap function (2.4) specialized as the feasibility gap (2.6) together with $D_{1}=0$, to obtain (3.6).
Moreover, under the step size condition (3.1) it follows that every weak accumulation point $\tilde{z}$ of the averaged trajectory $(\bar{z}^{k})$ satisfies
$\langle{\mathsf{F}}_{2}(z),\tilde{z}-z\rangle+g_{2}(\tilde{z})-g_{2}(z)\leq 0\hskip 17.00024pt\forall z\in\mathcal{U}_{2}$
In particular, $\Theta_{\rm Feas}(\tilde{z}|\mathcal{U}_{2})\leq 0$.
@@
Choosing constant step $t_{k}=t$ and penalty sequence as in (3.9), we easily see that $\frac{\sum_{k=1}^{K}t_{k}\sigma_{k}}{T_{K}}\leq\frac{a}{(1-\delta)(K+b)^{\delta}},$ so that the concrete complexity bound becomes
$\Theta_{\rm Feas}(\bar{z}^{K}|\mathcal{U}_{2})\leq\frac{\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{2Kt}+\frac{aC_{\mathcal{U}_{2}}}{(1-\delta)(K+b)^{\delta}}.$
@@
Corollary 3.2 is formulated for regularization sequences $(\sigma_{k})_{k\geq 1}$ of the form (3.9) with $\delta\in(0,1)$. In the limiting case $\delta=1$ and constant step size $t_{k}=t$, the complexity bound
$\Theta_{\rm Feas}(\bar{z}^{K}|\mathcal{U}_{2})\leq\frac{\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{2Kt}+\frac{a\log(K+b)}{K}C_{\mathcal{U}_{2}}.$
Hence, even in this limiting case, not satisfying condition (3.1) explicitly requested in Algorithm 1, the ergodic average $(\bar{z}^{k})_{k}$ asymptotically approaches the feasible set $\mathcal{S}_{2}$. $\Diamond$
@@
Let $\mathcal{U}_{1}\subset\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ be a compact set with $\mathcal{U}_{1}\cap\mathcal{S}_{1}\neq\varnothing$. Using the monotonicity of ${\mathsf{F}}_{1}$ and ${\mathsf{F}}_{2}$, we see
$\Psi_{k}(z)$ $=\langle{\mathsf{F}}_{2}(z^{k+1/2}),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z)$
$+\sigma_{k}\left(\langle{\mathsf{F}}_{1}(z^{k+1/2}),z^{k+1/2}-z\rangle+g_{1}(z^{k+1/2})-g_{1}(z)\right)$
$\geq\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z)$
$+\sigma_{k}\left(\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+g_{1}(z^{k+1/2})-g_{1}(z)\right).$
Hence, when choosing $z\in\mathcal{S}_{2}$, we are able to continue the above bound as
$\Psi_{k}(z)$ $\geq\sigma_{k}\left(\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+g_{1}(z^{k+1/2})-g_{1}(z)\right)\qquad\forall k\geq 1.$
Plugging this in (4.12) and dropping the non-positive term $\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2}$, we arrive at
$\sigma_{k}t_{k}\left(\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+g_{1}(z^{k+1/2})-g_{1}(z)\right)$ $\leq(E_{k}(z)+D_{k})-(E_{k+1}(z)+D_{k+1})$
$=W_{k}(z)-W_{k+1}(z).$
To continue with our estimation, we proceed with the following simple result.
Lemma 4.4.
Let either Assumption 3 hold or $\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ be compact. Then, the mapping $z\mapsto W_{k}(z)$ is bounded over bounded sets.
Proof.
Since
$W_{k}(z)$ $=\frac{1}{2}\norm{z^{k}-z}^{2}+\frac{t_{k}^{2}}{2}\norm{{\mathsf{V}}_{k}(z^{k+1/2})-{\mathsf{V}}_{k}(z^{k-1/2})}^{2}\leq\frac{1}{2}\norm{z^{k}-z}^{2}+\frac{t^{2}_{k}L^{2}_{k}}{2}\norm{z^{k+1/2}-z^{k-1/2}}^{2}$
the result follows immediately from the boundedness of the sequences $(z^{k})_{k\geq 1}$, $(z^{k-1/2})_{k\geq 1}$ that hold both under Assumption 3 and when $\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ is compact. $\blacksquare$
Since $\mathcal{S}_{1}\subseteq\mathcal{S}_{2}$, it follows from the assumptions made that $\mathcal{U}_{1}\cap\mathcal{S}_{2}\neq\varnothing$. Therefore, we can pick $z\in\mathcal{U}_{1}\cap\mathcal{S}_{2}$ arbitrary, so that we can find a positive constant $\bar{\omega}(z)$ such that $W_{k}(z)\leq\bar{\omega}(z)$ for all $k\geq 1$. Using this bound, and $\sigma_{k+1}\leq\sigma_{k}$, we arrive that
$\sum_{k=1}^{K}t_{k}\left(\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+g_{1}(z^{k+1/2})-g_{1}(z)\right)$ $\leq\frac{1}{\sigma_{1}}W_{1}(z)+W_{2}(z)(\frac{1}{\sigma_{2}}-\frac{1}{\sigma_{1}})+\ldots+W_{K}(z)(\frac{1}{\sigma_{K}}-\frac{1}{\sigma_{K-1}})$
@@
Dividing both sides by $T_{K}$ and applying the Jensen inequality, we obtain
$\langle{\mathsf{F}}_{1}(z),\bar{z}^{K}-z\rangle+g_{1}(\bar{z}^{K})-g_{1}(z)\leq\frac{\bar{\omega}(z)}{T_{K}\sigma_{K}}.$
Thus, for any compact subset $\mathcal{U}_{1}\subset\mathcal{Z}$ with $\mathcal{U}_{1}\cap\mathcal{S}_{2}\neq\varnothing$, we can use the definition of the localized gap function (2.4) specialized as the optimality gap (2.7) to obtain
$\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2})\leq\frac{C_{\mathcal{U}_{1}}}{T_{K}\sigma_{K}},$ (4.23)
where $C_{\mathcal{U}_{1}}\triangleq\sup_{z\in\mathcal{U}_{1}\cap\mathcal{S}_{2}}\bar{\omega}(z)$. This is the upper bound in (3.7). The lower bound in (3.7) holds since $\mathcal{S}_{1}\subseteq\mathcal{S}_{2}$:
@@
If additionally the lower level solution set enjoys $(\alpha,\rho)$ weak-sharpness we have from (2.9) and (3.6)
@@
$\geq-B_{\mathcal{U}_{1}}\left[\frac{\rho}{\alpha}\Theta_{\rm Feas}(\bar{z}^{K}|\mathcal{U}_{2})\cap\mathcal{S}_{2})\right]^{1/\rho}\geq-B_{\mathcal{U}_{1}}\left[\frac{\rho}{\alpha}\Theta_{\rm Feas}(\bar{z}^{K}|\mathcal{U}_{2})\right]^{1/\rho}$
@@
Choosing the penalty sequence $(\sigma_{k})_{k\geq 1}$ according to (3.9) and the step size $t_{k}=t$ small enough, the lower and upper complexity bound on the gap function becomes
$B_{\mathcal{U}_{1}}\left[\frac{\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{2Kt(\alpha/\rho)}+\frac{aC_{\mathcal{U}_{2}}}{(\alpha/\rho)(1-\delta)(K+b)^{\delta}}\right]^{1/\rho}\leq\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2})\leq\frac{C_{\mathcal{U}_{1}}}{at(K+b)^{1-\delta}}.$
Note that if $\delta=1$ in the choice of the penalty sequence (3.9), we do not obtain convergence in terms of the gap function to 0, but rather only an $O(1)$ upper bound and $o(1)$ lower bound.

### v2

@@ 4.1 Energetic bounds
We note that (B.2) in Appendix B guarantees that the upper bound is informative, in the sense that $c_{k}\geq 0$. Plugging this back into (4.16) and using (4.14), we obtain
$W_{k+1}(z^{\ast})\leq W_{k}(z^{\ast})-\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2}-\frac{1}{4}\norm{z^{k+1}-z^{k+1/2}}^{2}+t_{k}c_{k}.$
The summability condition (3.4) implies $\sum_{k}t_{k}c_{k}<\infty$. Hence, the non-negativity of $W_{k}(z^{*})$ allows us to deduce from Lemma A.1 that $\lim_{k\to\infty}W_{k}(z^{\ast})=W_{\infty}(z^{\ast})$ exists, and $\lim_{k\rightarrow\infty}\norm{z^{k+1/2}-z^{k}}=\lim_{k\to\infty}\norm{z^{k+1}-z^{k+1/2}}=0$. By the triangle inequality, the latter also implies that $\lim_{k\to\infty}\norm{z^{k+1/2}-z^{k-1/2}}=0$. By the definition of $E_{k}(z^{*})$, we further obtain that the sequences $(z^{k})_{k\geq 1}$ and $(z^{k+1/2})_{k\geq 1}$ are bounded.
By (4.3), $L_{k}$-Lipschitzness of ${\mathsf{V}}_{k}$ and item (b), we have, as $k\to\infty$,
$0\leq D_{k+1}$ $=t_{k}^{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}\leq t_{k}^{2}L^{2}_{k}\norm{z^{k+1/2}-z^{k-1/2}}^{2}\to 0.$
Since $E_{k}(z^{*})=W_{k}(z^{*})-D_{k}$, it thus follos $\lim_{k\rightarrow\infty}E_{k}(z^{*})=W_{\infty}(z^{*}).$ Moreover, from $\norm{z^{k}-z^{k+1/2}}\to 0$, and (4.1), we immediately deduce
$\lim_{k\rightarrow\infty}\norm{z^{k+1/2}-z^{*}}^{2}$ $=\lim_{k\rightarrow\infty}\norm{z^{k}-z^{*}}^{2}+\norm{z^{k}-z^{k+1/2}}^{2}+\langle z^{k}-z^{k+1/2},z^{k}-z^{*}\rangle$
$=\lim_{k\rightarrow\infty}\norm{z^{k}-z^{*}}^{2}=2W_{\infty}(z^{*}).$
This proves (d). $\blacksquare$
@@
Consider a point $z\in\mathcal{D}_{1}\cap\mathcal{D}_{2}$. Invoking Assumption 1, we have
$\langle{\mathsf{F}}_{1}(z^{k+1/2}),z-z^{k+1/2}\rangle\leq\norm{{\mathsf{F}}_{1}(z^{k+1/2})}\cdot\norm{z-z^{k+1/2}}$
$\leq\left(\norm{{\mathsf{F}}_{1}(z^{1})}+\norm{{\mathsf{F}}_{1}(z^{k+1/2})-{\mathsf{F}}_{1}(z^{1})}\right)\cdot\left(\norm{z^{1}-z^{k+1/2}}+\norm{z^{1}}+\norm{z}\right)$
$\leq\left(\norm{{\mathsf{F}}_{1}(z^{1})}+L_{{\mathsf{F}}_{1}}R\right)\cdot(R+\norm{z^{1}}+\norm{z})\triangleq a(z),$
@@
$g_{1}(z)-g_{1}(z^{k+1/2})$ $\leq\operatorname{\mathsf{Var}}(g_{1}|\{z\}\times\mathcal{B}_{R})\triangleq b(z).$
Thus, defining $C_{z}\triangleq a(z)+b(z)$ we can conclude that
$H^{({\mathsf{F}}_{1},g_{1})}(z,z^{k+1/2})=\langle{\mathsf{F}}_{1}(z^{k+1/2}),z-z^{k+1/2}\rangle+g_{1}(z)-g_{1}(z^{k+1/2})\leq C_{z}.$ (4.17)
By (4.15) and monotonicity of ${\mathsf{F}}_{2}$ we have, for all $z\in\mathcal{D}_{1}\cap\mathcal{D}_{2}$ and $k\geq 1$,
$\Psi_{k}(z)+\sigma_{k}C_{z}$ $\geq\Psi_{k}(z)+\sigma_{k}H^{({\mathsf{F}}_{1},g_{1})}(z,z^{k+1/2})$
$=-H^{({\mathsf{F}}_{2},g_{2})}(z,z^{k+1/2})\geq H^{({\mathsf{F}}_{2},g_{2})}(z^{k+1/2},z).$ (4.18)
Multiplying by $t_{k}$, using (4.14) and (4.5), where we now neglect the non-positive quadratic terms, we obtain, for all $z\in\mathcal{D}_{1}\cap\mathcal{D}_{2}$ and $k\geq 1$,
$t_{k}H^{({\mathsf{F}}_{2},g_{2})}(z^{k+1/2},z)$ $\leq t_{k}\Psi_{k}(z)+t_{k}\sigma_{k}C_{\mathcal{U}_{2}}\leq W_{k}(z)-W_{k+1}(z)+t_{k}\sigma_{k}C_{z}.$ (4.19)
Summing (4.19) for $k=1,\ldots,K$ dividing by $T_{K}$, and applying $D_{1}=0$, yields
$\sum_{k=1}^{K}\frac{t_{k}}{T_{K}}H^{({\mathsf{F}}_{2},g_{2})}(z^{k+1/2},z)\leq\frac{E_{1}(z)}{T_{K}}+\frac{\sum_{k=1}^{K}t_{k}\sigma_{k}}{T_{K}}C_{z}\qquad\forall z\in\mathcal{D}_{1}\cap\mathcal{D}_{2}.$ (4.20)
@@
$H^{({\mathsf{F}}_{2},g_{2})}(\bar{z}^{K},z)\leq\frac{E_{1}(z)}{T_{K}}+\frac{\sum_{k=1}^{K}t_{k}\sigma_{k}}{T_{K}}C_{z}\hskip 17.00024pt\forall z\in\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2}).$ (4.21)
Since $\mathcal{U}_{2}\subset\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ is compact, we have that $b(z)\leq\operatorname{\mathsf{Var}}(g_{1}|\mathcal{U}_{2}\times\mathcal{B}_{R})$ and $C_{\mathcal{U}_{2}}\triangleq\sup_{z\in\mathcal{U}_{2}}C_{z}$, as well as $\sup_{z\in\mathcal{U}_{2}}E_{1}(z)$, are well-defined. Thus, for any compact subset $\mathcal{U}_{2}\subset\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})\subset\mathcal{Z}$ we can use the definition of the localized gap function (2.4) specialized as the feasibility gap (2.7), to obtain (3.6).
@@
Choosing constant step $t_{k}=t$ and penalty sequence as in (3.9), we easily see that $\frac{\sum_{k=1}^{K}t_{k}\sigma_{k}}{T_{K}}\leq\frac{a}{(1-\delta)(K+b)^{\delta}},$ so that (3.6) transforms into (3.10).
@@
Corollary 3.2 is formulated for regularization sequences $(\sigma_{k})_{k\geq 1}$ of the form (3.9) with $\delta\in(0,1)$. In the limiting case $\delta=1$ and constant step size $t_{k}=t$, the last term in (3.6) is upper bounded as $\frac{a\log(K+b)}{K}C_{\mathcal{U}_{2}}$. Hence, even in this limiting case, not satisfying condition (3.1) required in Algorithm 1, the ergodic average $(\bar{z}^{k})_{k\geq 1}$ asymptotically approaches the feasible set $\mathcal{S}_{2}$. $\Diamond$
@@
Let $\mathcal{U}_{1}\subset\mathcal{D}_{1}\cap\mathcal{D}_{2}$ be a compact set s.t. $\mathcal{U}_{1}\cap\mathcal{S}_{1}\neq\varnothing$. By the monotonicity of ${\mathsf{F}}_{1}$ and ${\mathsf{F}}_{2}$ and choosing $z\in\mathcal{S}_{2}$, we see
[…]
$\geq\sigma_{k}H^{({\mathsf{F}}_{1},g_{1})}(z^{k+1/2},z)+H^{({\mathsf{F}}_{2},g_{2})}(z^{k+1/2},z)\geq\sigma_{k}H^{({\mathsf{F}}_{1},g_{1})}(z^{k+1/2},z).$
Plugging this in (4.5), dropping the non-positive quadratic terms, and recalling (4.14), we arrive at $\sigma_{k}t_{k}H^{({\mathsf{F}}_{1},g_{1})}(z^{k+1/2},z)\leq W_{k}(z)-W_{k+1}(z)$ for all $z\in\mathcal{S}_{2}$.
Since by the $L_{k-1}$-Lipschitz property of ${\mathsf{V}}_{k-1}$,
$W_{k}(z)$ $\leq\frac{1}{2}\norm{z^{k}-z}^{2}+t^{2}_{k-1}L^{2}_{k-1}\norm{z^{k-3/2}-z^{k-1/2}}^{2},$
we immediately see that, for any fixed $z\in\mathcal{S}_{2}$, $W_{k}(z)$ is bounded in the sense that for some positive constant $\bar{\omega}(z)$, $W_{k}(z)\leq\bar{\omega}(z)$ for all $k\geq 1$. Using this bound, $W_{k}(z)\geq 0$, and $\sigma_{k+1}\leq\sigma_{k}$, we obtain that, for any $z\in\mathcal{S}_{2}$,
$\sum_{k=1}^{K}t_{k}H^{({\mathsf{F}}_{1},g_{1})}(z^{k+1/2},z)$ $\leq\frac{1}{\sigma_{1}}W_{1}(z)+W_{2}(z)(\frac{1}{\sigma_{2}}-\frac{1}{\sigma_{1}})+\ldots+W_{K}(z)(\frac{1}{\sigma_{K}}-\frac{1}{\sigma_{K-1}})$
@@
Dividing both sides by $T_{K}$ and applying the Jensen inequality, we arrive at
$H^{({\mathsf{F}}_{1},g_{1})}(\bar{z}^{K},z)\leq\frac{\bar{\omega}(z)}{T_{K}\sigma_{K}}\qquad\forall z\in\mathcal{S}_{2}.$ (4.22)
Thus, for any compact subset $\mathcal{U}_{1}\subset\mathcal{Z}$ with $\mathcal{U}_{1}\cap\mathcal{S}_{2}\neq\varnothing$, we can use the definition of the localized gap function (2.4) specialized as the optimality gap (2.8) to obtain $\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2})\leq\frac{C_{\mathcal{U}_{1}}}{T_{K}\sigma_{K}}$, where $C_{\mathcal{U}_{1}}\triangleq\sup_{z\in\mathcal{U}_{1}\cap\mathcal{S}_{2}}\bar{\omega}(z)$. Note that the latter is well-defined since $\mathcal{S}_{1}\subseteq\mathcal{S}_{2}$ and $\mathcal{U}_{1}\cap\mathcal{S}_{1}\neq\varnothing$ by the Theorem assumption. Thus, we have obtained the upper bound in (3.7). The lower bound in (3.7) holds since $\mathcal{S}_{1}\subseteq\mathcal{S}_{2}$:
@@
If additionally the lower level solution set enjoys $(\alpha,\rho)$-weak sharpness we have from (2.10) and (3.6)
@@
$\geq-B_{\mathcal{U}_{1}}\left[\frac{\rho}{\alpha}\Theta_{\rm Feas}(\bar{z}^{K}|\mathcal{U}_{2}\cap\mathcal{S}_{2})\right]^{1/\rho}\geq-B_{\mathcal{U}_{1}}\left[\frac{\rho}{\alpha}\Theta_{\rm Feas}(\bar{z}^{K}|\mathcal{U}_{2})\right]^{1/\rho}$
@@
Choosing the penalty sequence $(\sigma_{k})_{k\geq 1}$ according to (3.9) and the step size $t_{k}=t$, the r.h.s. of (3.7) is upper bounded by the r.h.s. in (3.11) and the l.h.s. in (3.11) follows from (3.10). Note that if $\delta=1$ in the choice of the penalty sequence (3.9), we do not obtain convergence in terms of the gap function to 0, but rather only an $O(1)$ upper bound and a $o(1)$ lower bound.
Proof of Theorem 3.3
(i) From the prox-inequality (2.1) applied to the point $z^{k+1/2}$ and divided by $t_{k}>0$, we have, for any $z\in\mathcal{Z}$,
$t_{k}^{-1}\langle z^{k+1/2}-z^{k},z^{k+1/2}-z\rangle$ $\leq\langle{\mathsf{V}}_{k}(z^{k-1/2}),z-z^{k+1/2}\rangle+G_{k}(z)-G_{k}(z^{k+1/2})$
[…]
$\hskip-17.00024pt\leq\sigma_{k}H^{({\mathsf{F}}_{1},g_{1})}(z,z^{k+1/2})-H^{({\mathsf{F}}_{2},g_{2})}(z^{k+1/2},z)+e_{k},$ (4.23)
[…]
$\lim_{j\to\infty}H^{({\mathsf{F}}_{1},g_{1})}(z,z^{k_{j}+1/2})=\lim_{j\to\infty}\langle{\mathsf{F}}_{1}(z^{k_{j}+1/2}),z-z^{k_{j}+1/2}\rangle+g_{1}(z)-g_{1}(z^{k_{j}+1/2})$
[…]
Thus, since $\sigma_{k}\to 0$ as $k\to\infty$, we have for any $z\in\operatorname{dom}(g_{2})\subseteq\operatorname{dom}(g_{1})$, that $\lim_{j\to\infty}\sigma_{k_{j}}H^{({\mathsf{F}}_{1},g_{1})}(z,z^{k_{j}+1/2})\leq 0.$ Thus, we obtained the limits of all the terms in (4.23) except the term with $H^{({\mathsf{F}}_{2},g_{2})}$. Using these limits when passing in (4.23) to the subsequence indexed by $k_{j}$ and then to the limit $j\to\infty$, using the lower semicontinuity of $g_{2}$, we obtain, for any $z\in\operatorname{dom}(g_{2})$
$0$ $\leq\lim_{j\to\infty}-H^{({\mathsf{F}}_{2},g_{2})}(z^{k_{j}+1/2},z)=\lim_{j\to\infty}\langle{\mathsf{F}}_{2}(z),z-z^{k_{j}+1/2}\rangle+g_{2}(z)-g_{2}(z^{k_{j}+1/2})$
[…]
For $z\notin\operatorname{dom}(g_{2})$ the above inequality holds trivially. Thus, it holds for all $z\in\mathcal{Z}$. This shows that any weak limit point of $(z^{k+1/2})_{k\geq 1}$ is a Minty solution of $\operatorname{HVI}({\mathsf{F}}_{2},g_{2})$. Under Assumption 1, Minty solutions and Stampacchia solutions of $\operatorname{HVI}({\mathsf{F}}_{2},g_{2})$ (2.3) coincide (see also the proof of Lemma 2.1 in Appendix C), showing that $\tilde{z}\in\mathcal{S}_{2}$.
(ii) By Lemma 4.2 (c), the sequence $(z^{k+1/2})_{k\geq 1}$ is bounded. Hence the ergodic average sequence $(\bar{z}^{k})_{k\geq 1}$ is bounded as well and thus admits weakly converging subsequences. Let $\tilde{z}$ be such that $\bar{z}^{k_{j}}\rightharpoonup\tilde{z}$ as $j\to\infty$. For any reference point $z\in\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})=\operatorname{dom}(g_{2})$ applying condition (3.1) to (4.21), we obtain that $\lim_{j\rightarrow\infty}H^{({\mathsf{F}}_{2},g_{2})}(\bar{z}^{k_{j}},z)\leq 0$. Thus, using the lower semi-continuity of $g_{2}$, $\tilde{z}$ satisfies $H^{({\mathsf{F}}_{2},g_{2})}(\tilde{z},z)\leq 0$. This inequality trivially holds for $z\notin\operatorname{dom}(g_{2})$, and hence holds for all $z\in\mathcal{Z}$. Thus, $\tilde{z}$ is a Minty solution of $\operatorname{HVI}({\mathsf{F}}_{2},g_{2})$, which again, under Assumption 1 is also a Stampacchia solution of $\operatorname{HVI}({\mathsf{F}}_{2},g_{2})$. This implies $\tilde{z}\in\mathcal{S}_{2}$.
For any point $z\in\mathcal{S}_{2}$, (4.22) and (3.1) yield that $\lim_{j\rightarrow\infty}H^{({\mathsf{F}}_{1},g_{1})}(\bar{z}^{k_{j}},z)\leq 0.$ This, by the lower semi-continuity of $g_{1}$, implies
$\langle{\mathsf{F}}_{1}(z),\tilde{z}-z\rangle+$ $g_{1}(\tilde{z})-g_{1}(z)\leq\lim_{j\to\infty}\left(\langle{\mathsf{F}}_{1}(z),\bar{z}^{k_{j}}-z\rangle+g_{1}(\bar{z}^{k_{j}})-g_{1}(z)\right)\leq 0.$
Since $z\in\mathcal{S}_{2}$ is arbitrary and $\tilde{z}\in\mathcal{S}_{2}$, we have that $\tilde{z}$ is a Minty solution of (P), which is, under Assumption 1, also a Stampacchia solution of $\operatorname{HVI}({\mathsf{F}}_{1},g_{1})$, i.e., $\tilde{z}\in\mathcal{S}_{1}$.
Lemma 4.2 (d) shows that the limit of $(\norm{z^{k+1/2}-z})_{k\geq 1}$ exists for all $z\in\mathcal{S}_{1}$. As just demonstrated, all weak accumulation points of $(\bar{z}^{k})_{k\geq 1}$ are points in $\mathcal{S}_{1}$. Lemma A.2 thus shows that $\bar{z}^{k}\rightharpoonup z_{\infty}$ for some $z_{\infty}\in\mathcal{S}_{1}$.